OpenStudy (anonymous):

could someone help me set this up i have not done one of these problems in a while : And on his farm he had some cows and chicken. He counted a total of 123 heads and 456 legs. How many cows does Old McDonald have?

OpenStudy (johnweldon1993):

Well you know cows have 4 legs and chickens have 2 legs.... And you know that both of them only have 1 head lol ...so that means let cow = c and chicken = n c + n = 123 **each one can only have 1 head...so yeah* and we also know that since cows have 4 legs and chickens have 2...and there are a total of 456 legs...we have 4c + 2n = 456 a system of equations c + n = 123 4c + 2n = 456 can you solve it?

OpenStudy (anonymous):

yes thank you for your help :)

OpenStudy (johnweldon1993):

No problem! :)